Providence Hospital generates monthly performance reports for each of its departments. The hospital must maintain an adequate staff of attending and on-call physicians at all times, so physician costs are not affected by the number of patient visits. But all other costs do vary with patient activity. Nurse-hours are used as the activity measure for nursing costs, and patient visits are used as the activity measure for the cost of supplies and other variable costs.

The head physician of the hospital's emergency room, Yolanda Mortensen, is responsible for control of costs. During October, the emergency room unit expected to treat 3,900 patients but actually treated 3,600 patients. The following additional information for October is available:

Budget

Actual

Variance

Nurse-hours

2,340

2,232

108

Nursing costs

$51,480

$30,100

$21,380

Supplies & other variable costs

$39,000

$29,000

$10,000

Fixed costs

$106,300

$96,400

$9,900


Required
Compute the flexible-budget variances for each of the cost categories for October (NOTE: enter favorable variances as positive numbers and unfavorable variances as negative numbers):

  Nursing costs    

  Supplies and other variable costs    

  Fixed costs   
Answer #1
Given,
Budget Actual
Nurse-hours 2340 2232
Nursing cost 51480 30100
Supplies and other variable costs 39000 29000
Fixed costs 106300 96400
Standard nursing cost/hour = 51,480/2,340 = 22
Budgeted number of patient visits = 3,900
Supplies and other variable costs / patient visit = 39,000/3,900 = 10
Static Budget Actual Variance
Nursing cost 2,232 * 22
= 49,104
         30,100          19,004
Supplies and other variable costs 3,600 * 10
= 36,000
         29,000            7,000
Fixed costs            106,300          96,400            9,900